1.) Rina needs at least 48 units of protein, 60 units of carbohydrates, and 50 units of fat each month. From each kilogram of food A, she receives 2 units of protein, 4 units of carbohydrates, and 5 units of fats. Food b contains 3 units of protein, 3 units of carbohydrates, and 2 units of fats. If food A costs Php110 per kilogram and food B costs Php 90 per kilogram. How many kilograms of each food should Rina buy each month to keep costs at a minimum?


1
Expert's answer
2022-01-11T00:59:50-0500

minimize costs:

"z=110x_1+90x_2"

subject to:

"2x_1+3x_2\ge 48"

"4x_1+3x_2\ge 60"

"5x_1+2x_2\ge 50"

x1, x2 are kg of foods A and B




for Extreme Points:

for "5x_1+2x_2= 50" :

"x_1=0\implies x_2=25"

for "2x_1+3x_2= 48" :

"x_2=0\implies x_1=24"

intersection "5x_1+2x_2= 50" and "4x_1+3x_2= 60" :

"x_1=4.29,x_2=14.29"

intersection "2x_1+3x_2= 48" and "4x_1+3x_2= 60" :

"x_1=6,x_2=12"


Objective function values at Extreme Points:

"z(0,25)=2250"

"z(4.29,14.29)=1757.14"

"z(6,12)=1740"

"z(24,0)=2640"


The miniimum value of the objective function z=1740 occurs at the extreme point (6,12).

Hence, the optimal solution to the given LP problem is : x1=6, x2=12 and min z=1740.